OpenStudy (anonymous):

if possible, evaluate f(x) at the given values of x f(x)=x1/4 x=81,-2 f(81)=? f(-2)=?

OpenStudy (anonymous):

Just do 81 * (1/4) and -2 * (1/4)

OpenStudy (anonymous):

how do I multiply 81 times 1/4 is it 325?

OpenStudy (anonymous):

81 * (1 / 4) is the same as 81 / 4. X * (Y / Z) = XY/Z

OpenStudy (anonymous):

you're thinking of X = Y/Z which is XZ = Y.

OpenStudy (anonymous):

i will be willing to bet that it is \[f(x)=x^{\frac{1}{4}}\]

OpenStudy (anonymous):

in which case \[f(81)=3\]

OpenStudy (anonymous):

what is f(-2)

OpenStudy (anonymous):

if my guess is right, then \[f(-2)=\sqrt[4]{-2}\] which is not a real number

OpenStudy (anonymous):

thank u.....

OpenStudy (anonymous):

so IF i am correct the domain of \[f(x)=\sqrt[4]{x}\] had domain \[x\geq0\]

OpenStudy (anonymous):

thank u...